OpenStudy (anonymous):

What is the integral of sin(8x)cos(8x)^(1/3)dx?

OpenStudy (anonymous):

It looks like this in equation form: \[\int\limits_{}^{} \sin(8x)\sqrt[3]{\cos(8x)}dx\]

OpenStudy (anonymous):

My answer was \[-6\cos ^{\frac{ 4 }{ 3 }}(8x)+c\] and i tried it twice but couldn't find my mistake...

OpenStudy (raden):

integral by u-sub let u = cos(8x) du = -8sin(8x) dx -1/8 du = sin(8x) dx see, the integral becomes int (-1/8 u^(1/3)) du = - 1/8 int u^1/3 du = .... (this should be easy for you now)

OpenStudy (anonymous):

omg that was such a small mistake, i had -1/8 as my coefficient on my du instead of -8 -.- thanks for pointing that out

OpenStudy (raden):

you're welcome :D